OpenStudy (anonymous):

James owes $975 in 3.5 years and $1500 in 8.5 years. However, James is unable to meet his $975 obligation at the end of 3.5 years. By mutual agreement with the lender, James is allowed to pay off both obligations at the end of 6.5 years based on a simple interest rate of 10.75%. Determine the amount required to settle the debts at the end of 6.5 years.

OpenStudy (phi):

Do you know the formula for simple interest? http://math.about.com/od/businessmath/ss/Interest.htm

OpenStudy (anonymous):

Yes. I've done each formula for the 2 separate problems then added them together and thats not right.

OpenStudy (phi):

can you say exactly what you did?

OpenStudy (anonymous):

975(3.5)(.1075)= 366.84 1500(8.5)(.1075)= 1370.625 366.84+1370.625= 1737.465

OpenStudy (anonymous):

I also tried (975+1500)=2475+ (6.5+.1075)=4204.41

OpenStudy (phi):

They changed the original loan agreement pay off both obligations at the end of 6.5 years so where you have 975(3.5)(.1075)= 366.84 1500(8.5)(.1075)= 1370.625 you should use 6.5 years for both loans. then add up to get the total interest. Then add in the principal, because he has to pay pack the principal plus the interest. you said I also tried (975+1500)=2475+ (6.5+.1075)=4204.41 A couple things: the equation multiplies, not adds, and the numbers don't come out as 4204.41. But in any case, it should be Interest= P*time*rate, or 2475*6.5*0.1075= 1729.41 (which should match doing the 2 loans separately and adding the interest together) now add in the amount of the loan 2475 to get the final answer

OpenStudy (anonymous):

I did that and I got it wrong. Thanks for trying to help!

OpenStudy (anonymous):

the answer is $2524.01 fyi

OpenStudy (phi):

Looks like you are the victim of a bad question/answer If you owe a total of 2475 and after 6.5 years you pay 2524.01, that means you paid 49.01 in interest, or 7.54 per year. 7.54 per year on 2475 is not 10.75% interest. It is 0.3% . So something is definitely wrong with this question. I would take this one to your teacher, and send e-mail to whoever came up with it.